OpenStudy (ellerai):

what is a slope of a line perpendicular to the line with equation y=4x+5

Vocaloid (vocaloid):

any ideas?

OpenStudy (ellerai):

would i have to solve the equation? i don't know how to do this :(

Vocaloid (vocaloid):

no, no need to solve do you know how to find the slope of y = 4x + 5?

OpenStudy (ellerai):

slope intercept form?

Vocaloid (vocaloid):

good, but it's already in slope intercept form

OpenStudy (ellerai):

so the slope is 4

Vocaloid (vocaloid):

good! now, to find the perpendicular slope, you turn 4 into 1/4 (by taking the reciprocal) and multiply it by -1 to get -1/4

Vocaloid (vocaloid):

so -1/4 is your answer

OpenStudy (ellerai):

would you do the same for parallel?

Vocaloid (vocaloid):

for parallel, you just leave the slope alone

Vocaloid (vocaloid):

(since parallel lines have equal slopes)

Vocaloid (vocaloid):

so for parallel it would be 4

OpenStudy (ellerai):

what if it is not in slope intercept form?

Vocaloid (vocaloid):

change it into intercept form

Vocaloid (vocaloid):

would you like an example?

OpenStudy (ellerai):

yes please

Vocaloid (vocaloid):

ok, say we have -2x + y = 3

Vocaloid (vocaloid):

do you know how we could change it to y = mx + b form?

OpenStudy (ellerai):

by isolating y?

Vocaloid (vocaloid):

yes!

Vocaloid (vocaloid):

would you like to try giving it a go?

OpenStudy (ellerai):

yeah gimme a sec

OpenStudy (ellerai):

so i subtracted 2x from both sides and i got y = 3 - 2x

OpenStudy (ellerai):

@Vocaloid

Vocaloid (vocaloid):

almost

Vocaloid (vocaloid):

instead of subtracting, you actually want to add 2x instead

OpenStudy (ellerai):

but 2x wasn't negative?

Vocaloid (vocaloid):

since we had -2x initially, we must add 2x to get 0

Vocaloid (vocaloid):

our original problem was - 2x + y = 3 so yes, x was negative

OpenStudy (ellerai):

oh i didn't see the negative sorry

Vocaloid (vocaloid):

nah it's cool

Vocaloid (vocaloid):

so that gives us y = 3 + 2x. we can just switch the 3 and 2x to get y = 2x + 3

Vocaloid (vocaloid):

making our slope 2 (parallel) and -1/2 (perpendicular)

OpenStudy (ellerai):

oh okay!

OpenStudy (ellerai):

thank u!